# Composed Multi-modal Retrieval: A Survey of Approaches and Applications

### v1

@@ 4.1.2 Implicit visual Transformation
Existing works on implicit visual transformation for zero-shot composed image retrieval (ZS-CIR) can be categorized into three main groups: single pseudo-word methods [70, 16, 71, 72], multiple pseudo-word methods [73, 74], and other methods [75, 76].
@@
Single pseudo-word methods focus on mapping an image into a single pseudo-word token for textual representation. Saito et al. [70] propose Pic2Word, which maps CLIP visual features to a single pseudo-word token via a cycle contrastive loss, showing the powerful generalization across various ZS-CIR tasks using only weakly-supervised or unsupervised data. Similarly, Baldrati et al. [16] introduce SEARLE, an optimization-based textual inversion to generate pseudoword tokens and use knowledge distillation to train a Textual Inversion Network. SEARLE achieves efficient inference and improves performance. Moreover, they introduce an open-domain benchmarking dataset, CIRCO, which is the first CIR dataset to contain multiple ground-truth labels for each query. Expanding SEARLE, the iSEARLE proposed by Agnolucci et al. [71] further improves the retrieval accuracy by improving robustness with Gaussian noise, adding regularization to ensure dense token representations, and incorporating hard negative sampling to capture fine-grained details. Lastly, Context-I2W [72] employs a context-dependent mapping strategy, combining an Intent View Selector for dynamic view selection and a Visual Target Extractor for contextually relevant pseudo-word tokens, achieving significant performance gains.
@@
Multiple pseudo-word methods extend representation by mapping images to multiple pseudo-word tokens to better capture fine-grained details. Lin et al. [73] introduces fine-grained textual inversion network, FTI4CIR, where an image is represented by a subject-oriented pseudo-word token and several attribute-oriented pseudo-word tokens, aligned with real-word tokens using semantic regularization. This approach enhances the expressiveness of image content. Similarly, Du et al. [74] propose an asymmetric ZS-CIR framework ISA, which employs an adaptive token learner to map images into sentence-like representations with discriminative visual information, and combines global contrastive distillation and local alignment regularization to improve retrieval accuracy and efficiency, particularly in resource-constrained environments.
Some approaches focus on improving the integration of image and text representations or adapting pre-trained models for ZS-CIR. The Slerp proposed by Jang et al. [76] employs Spherical Linear Interpolation for direct merging of image and text representations and Text-Anchored Tuning (TAT) to refine the image encoder, narrowing the modality gap and preserving image representation integrity. Chen et al. [75] propose Masked Tuning to bridge the gap between pre-trained models and ZS-CIR tasks by masking image patches to generate triplets (masked image, text, image), and learning fine-grained text-guided modifications that capture more detailed information for improved retrieval.
While single pseudo-word methods prioritize efficient and robust representations, multiple pseudo-word methods enhance the granularity of visual-text alignment. Other methods innovate in merging modalities or adapting pre-trained models for ZS-CIR, showcasing diverse yet complementary strategies to advance the field.
@@ 4.2 Text-side Transformation
The ZS-CIR methods for image-side transformation rely on pre-defined simple prompts (e.g., “a photo of [\$]”), which leads to a very homogeneous input to text encoders and cannot fully learn the subtleties of natural language variations. This limitation becomes more significant especially when trying to scale up the model, as larger models require more diverse data to capture more complex semantic information.
@@
Gu et al. [77] and Byun et al. [78] focus on optimizing text-side transformations for CIR by addressing limitations in text encoding and modality alignment. The LinCIR proposed by Gu et al. leverages a self-supervised mechanism called Self-Masking Projection (SMP) to reduce dependency on triplet datasets and minimize modality gaps. This involves generating augmented texts through keyword masking and adding noise to textual embeddings, ensuring consistency between original and augmented representations. In contrast, Byun et al. introduce RTD, a plug-and-play training scheme that enhances text encoders through target-anchored contrastive learning, refined batch sampling with hard negatives, and a concatenation scheme, effectively addressing task discrepancies without requiring additional fine-tuning data.
Thawakar et al. [79] and Li et al. [80] emphasize integrating multimodal contexts to improve CIR and related tasks. Thawakar et al. propose CoVR, which uses rich language descriptions to encode query-specific contexts and learns discriminative visual-textual embeddings for accurate video retrieval, achieving sota results. Li et al. focus on multimodal compositional learning (MCL) for large language models, introducing tasks like Multimodal-Context Captioning and Retrieval to guide frozen language models in synthesizing multimodal information. Both methods enhance alignment and retrieval accuracy by leveraging complementary multimodal features, with MCL targeting more complex multimodal contexts, such as text generation and scene understanding.
4.3 External Knowledge Assistance
Suo et al. [81] introduce the Knowledge-Enhanced Dual-stream ZS-CIR (KEDs) framework to address the limitation of existing methods that overlook fine-grained attribute information. While traditional approaches mainly focus on global visual features, KEDs enhances pseudo-word tokens by integrating external knowledge from a database, which captures detailed attributes like color, object count, and layout. The framework includes two key components: a Bi-modal Knowledge-guided Projection (BKP) network that enriches pseudo-word tokens by retrieving relevant images and captions, and a Dual-stream Semantic Alignment module that explicitly aligns these enriched tokens with textual semantics. This approach leverages external knowledge to significantly improve the accuracy and domain-specific performance of ZS-CIR.
@@ 5 Semi-Supervised Learning-based CMR
In response to the problem that CIR datasets are small and difficult to scale, some works focus on semi-supervised learning-based Cross-Modal Retrieval (CMR), which can be broadly classified into four categories: (1) Automated data construction [1, 7]; (2) Synthetic data generation and dataset expansion [9, 5]; (3) Few-shot and semi-supervised learning [2, 6]; (4) Multimodal representation and retrieval enhancement [10, 8, 80]. Below, we will introduce the details of each category.
(1) Automated data construction: Liu et al. [1] introduce an automated data construction pipeline that leverages large-scale image-caption data to build training datasets for CIR task. The proposed TransAgg utilizes a transformer-based aggregation technique to dynamically fuse information from different modalities, demonstrating strong performance, particularly in zero-shot settings. Similarly, Levy et al. [7] focuse on overcoming the limitations of small datasets and modality redundancy in CIR systems. They introduce the LaSCo dataset, which is ten times larger than previous ones, and propose a CASE model, which integrates early fusion of modalities through a cross-attention mechanism. CASE improves retrieval performance on standard test sets. Both approaches highlight the critical role of large-scale, high-quality datasets and efficient fusion methods for enhancing CMR performance.
(2) Synthetic data generation and dataset expansion: Gu et al.[9] propose CompoDiff, a diffusion-based model designed to generate synthetic data for ZS-CIR. By creating a large-scale synthetic dataset, SynthTriplets18M, they extend the available training data, which significantly enhance the generalization ability of CIR models. CompoDiff also offers flexibility by handling conditions beyond traditional text, such as negative text and image mask conditions. Later, Jiang et al. [5] develope HyCIR, which uses synthetic labels generated from unlabeled images via a synthesis pipeline. This pipeline utilizes VLMs and LLMs to generate query text for each image pair, followed by semantic filtering. HyCIR employs a hybrid training strategy that combines zero-shot CIR supervision with synthetic triples to improve the generalization ability.
(3) Few-shot and semi-supervised learning: Hou et al. [2] address the challenge with limited labeled data by introducing a two-stage pseudo triplet guided few-shot CIR scheme (PTG-FSCIR). The first stage employs a masking training strategy and advanced image caption generator to construct pseudo triplets from pure image data; the second stage selects challenging unlabeled samples for fine-tuning based on active learning. This method demonstrates that carefully selected pseudo-data can significantly improve model performance, even with a small number of annotated samples. Jang et al. [6] propose Visual Delta Generator (VDG), a semi-supervised method that generates pseudo-triplets by describing visual differences between reference and target images using LLMs. The advantage of this method is that it can fully utilize existing large-scale image-text pair datasets without additional annotation costs.
(4) Multimodal representation and retrieval enhancement: Zhou et al. [10] propose VISTA, an embedding model for universal multimodal retrieval that overcomes the limitations of text-oriented retrievers by handling image-only and text-only data. VISTA generates high-quality composed image-text pairs using a template-based method and an LLM-based approach for diverse image descriptions. Then, VISTA employs a multi-stage training algorithm, aligning visual and textual encoders with weakly labeled data, and enhancing multimodal representation capability. Zhang et al. [8] introduce the CTI-IR model, which combines GANs and automatically manipulates the visual features of the reference image to match the textual description by jointly training the generative and retrieval models. Li et al. [80] introduce Multimodal Composition Learning (MCL), designed to improve LLMs’ understanding of complex visual and language contexts. MCL includes two tasks—Multimodal-Context Captioning and Multimodal-Context Retrieval—that enhance LLMs’ ability to process multimodal data. MCL shows strong performance on various benchmarks, especially in the zero-shot setting of the CIR task, revealing the potential of task-specific guidance for LLMs.
By leveraging techniques like automated data construction, synthetic data generation, few-shot learning, and advanced multimodal representations, these methods allow models to perform well and achieve high accuracy, even with limited data. These advancements improve model robustness and offer new solutions to the data bottleneck in CIR tasks.

### v2

@@ 4.1.2 Implicit Visual Transformation
Existing works on implicit visual transformation for zero-shot CMR can be categorized into three main groups: single pseudo-word methods [12, 23, 77, 78, 138, 139, 141], multiple pseudo-word methods [79, 80, 81, 140].
@@
Single pseudo-word methods focus on mapping an image into a single pseudo-word token for textual representation. Saito et al. [12] propose Pic2Word, which maps CLIP visual features to a single pseudo-word token via a cycle contrastive loss, showing the powerful generalization across various ZS-CMR tasks using only image-text data. Similarly, Baldrati et al. [23] introduce SEARLE, an optimization-based textual inversion to generate pseudo-word tokens and use knowledge distillation to train a Textual Inversion Network. SEARLE achieves efficient inference and improves performance. Expanding SEARLE, the iSEARLE proposed by Agnolucci et al. [77] further improves the retrieval accuracy by improving robustness with Gaussian noise, adding regularization to ensure dense token representations, and incorporating hard negative sampling to capture fine-grained details. Then, Context-I2W [78] employs a context-dependent mapping strategy, combining an Intent View Selector for dynamic view selection and a Visual Target Extractor for contextually relevant pseudo-word tokens generation, achieving significant performance gains. On this basis, Tang et al. [141] further propose Denoise-I2W, which integrates a learnable attention-based denoising module to suppress intention-irrelevant visual regions during image-to-word mapping, and leverages contrastive learning to enhance semantic consistency, achieving notable performance gains on multiple CIR benchmarks. Wang et al. [139] propose a framework, TSCIR, that generates pseudo-word tokens by modeling complex semantic structures such as attribute-object and action-target pairs through a structured text encoder and semantic alignment loss, improving performance in fine-grained ZS-CMR tasks. Lastly, Tang et al. [138] propose MTIP-WM, which predicts missing target-relevant semantics using a world model trained on pseudo-triplets, generating enriched pseudo-word tokens that capture both observed and inferred attributes, and then combines these tokens with manipulation text for robust contrastive learning.
@@
Multiple pseudo-word methods extend representation by mapping images to multiple pseudo-word tokens to better capture fine-grained details. Lin et al. [79] introduce a fine-grained textual inversion network, FTI4CIR, where an image is represented by a subject-oriented pseudo-word token and several attribute-oriented pseudo-word tokens, aligned with real-word tokens using semantic regularization. This approach enhances the expressiveness of image content. Similarly, Du et al. [80] propose an asymmetric ZS-CMR framework ISA, which employs an adaptive token learner to map images into sentence-like representations with discriminative visual information, and combines global contrastive distillation and local alignment regularization to improve retrieval accuracy and efficiency, particularly in resource-constrained environments. Suo et al. [81] introduce the Knowledge-Enhanced Dual-stream ZS-CMR (KEDs) framework to address the limitation of existing methods that overlook fine-grained attribute information. KEDs enhances pseudo-word tokens by integrating external knowledge from a database, which captures detailed attributes like color, object count, and layout. This approach leverages external knowledge to significantly improve the accuracy and domain-specific performance of ZS-CMR. Chen et al. [140] introduce a Data-efficient Generalization (DeG) framework for ZS-CMR, addressing modality and distribution shift challenges. The framework comprises two key components: the Textual Supplement (TS) module, which refines pseudo-word tokens by enhancing their semantic representation, and the Semantic Set (S-Set), which mitigates overfitting by leveraging the zero-shot capabilities of VLMs, improving the generalization. Recently, Li et al. [li2025rethinking] propose MOA, a object-aware pseudo-word learning framework that identifies query-relevant objects through a multi-object recognizer and a noun-guided filtering strategy, and then maps the screened objects into multiple pseudo-word tokens for precise ZS-CMsR. This simple yet effective design achieves competitive results on multiple benchmarks, significantly outperforming prior methods.
@@ 4.2 Text-side Transformation
The ZS-CMR methods for image-side transformation rely on pre-defined simple prompts (e.g., “a photo of [\$]”), which leads to a very homogeneous input to text encoders and cannot fully learn the subtleties of natural language variations. This limitation becomes more significant, especially when trying to scale up the model, as larger models require more diverse data to capture more complex semantic information.
Figure 5: Text-side Transformation. Figure is from [82].
@@
Recently, text-side transformations for CMR have been proposed to address the above limitations. As shown in Fig. 5, the LinCIR proposed by Gu et al. [82] leverages a self-supervised mechanism called self-masking projection to reduce dependency on triplet datasets and minimize modality gaps. This involves generating augmented texts through keyword masking and adding noise to textual embeddings, ensuring consistency between original and augmented representations. Building on this, Li et al. [86] propose MoTaDual, a multimodal-task dual alignment framework that uses LLMs (e.g., LLaMA, GPT-4) to generate diverse modification instructions and target texts, and applies prompt tuning to a lightweight textual inversion network, effectively addressing both modality and task discrepancies and achieving strong performance across four ZS-CMR benchmarks. Other methods also propose text-side enhancement strategies. Byun et al. [83] introduce RTD, a plug-and-play training scheme that enhances text encoders through target-anchored contrastive learning, refined batch sampling with hard negatives, and a concatenation scheme, effectively addressing task discrepancies without requiring additional fine-tuning data. Thawakar et al. [84] propose CoVR, which uses rich language descriptions to encode query-specific contexts and learns discriminative visual-textual embeddings for accurate video retrieval, achieving SOTA results. Li et al. [85] focus on multimodal compositional learning for LLMs, introducing tasks like Multimodal-Context Captioning and Retrieval to guide frozen language models in synthesizing multimodal information.
4.3 Linear Interpolation
As shown in Fig. 6, Other approaches focus on improving the integration of image and text representations for ZS-CMR. Early work ALIGN proposed by Jia et al. [87] attempts to directly use the linear image+text features as multimodal queries to retrieve the intended images. The Slerp proposed by Jang et al. [88] employs Spherical Linear Interpolation for direct merging of image and text representations and Text-Anchored Tuning (TAT) to refine the image encoder, narrowing the modality gap and preserving image representation integrity. Jang et al. further [90] propose Slerp++, which introduces semantics-aware dynamic interpolation weights to adaptively control the fusion of image and text features in high-dimensional space, and combines it with text-anchored tuning to keep language representations stable, enabling more accurate composed embeddings for ZS-CMR. Chen et al. [89] propose Masked Tuning to bridge the gap between pre-trained models and ZS-CMR tasks by masking image patches to generate triplets (masked image, text, image). During inference, the model uses a linearly weighted combination of image and text features as a multimodal query, where the weight is a predefined mask rate to alleviate the distribution shift between masked images in pre-training and complete images in inference.
Figure 6: Linear Interpolation for ZS-CMR. Figure is from [88].
@@ 5 Semi-Supervised Learning-based CMR
Early supervised works primarily relied on manual annotation to construct triplet datasets. While these efforts provided a foundation for dataset construction, the manual annotation process is time-consuming and labor-intensive, leading to several significant limitations: (1) The data scale and domains are limited, constrained by the human effort required for annotation, impairing generalizability; (2) The modified text is frequently simplistic, lacking the complexity required for real-world applications; (3) Annotations may have noise and uncertainty, which becomes a bottleneck for further advancements in the CMR field.
To mitigate these issues, the semi-supervised learning-based CMR (SSL-CMR) paradigm is proposed, which generates pseudo-triplets via image and text generation techniques. These approaches aim to improve the quantity of triplet data while having high-quality triplet annotations by minimizing the introduction of noise and uncertainty. These advancements allow SSL-CMR models to perform well and achieve high accuracy, improving model robustness and offering new solutions to the data bottleneck in CMR tasks. In the following, we highlight related key works and their technical innovations.
5.1 Automatic Data Construction
Works in this research line can be divided into two major categories: (1) generating language content only [91, 92, 71, 93, 94, 95, 96, 146]; and (2) generating language and visual content simultaneously [97, 98, 73, 111].
5.1.1 Generating Language Content
The augmentation of triplet data through text generation can be categorized into two main approaches: (1) generating text from a single image and (2) generating text describing differences between two images.
In the first category, Liu et al. [91] propose a pipeline leveraging large-scale image-text datasets (e.g., Laion-COCO) to synthesize relative descriptions. They use predefined templates and LLMs such as ChatGPT to edit image captions, forming triplets by matching generated text with candidate image captions. This yields two datasets: Laion-CIR-Template and Laion-CIR-LLM, each with approximately 16K triplets. Template-based descriptions demonstrate superior performance in fine-grained retrieval tasks. Hou et al. [92] introduce a masked training method, where patches of the original image are masked to create a reference image. Captions generated for the original (target) image are used as modification text, forming pseudo triplets. Zhang et al. [71] focus on generating challenging negative examples. They modified query sentences or replaced reference images with semantically mismatched alternatives to create pseudo triplets with subtle differences. This improved the model’s ability to distinguish examples and learn a better metric space for complex retrieval tasks. Liu et al. [93] introduce bidirectional training, where the model is trained on both forward (reference image + modification text to target image) and backward (target image + reverse modification text to reference image) queries. This expands data volume and strengthens multimodal fusion capabilities. Jawade et al. [144] propose SCOT, which generates pseudo-triplets by masking parts of the image to create reference images, with captions for the target images used as modification text. These pseudo-triplets are then used for training, where a self-training process with a teacher-student framework is applied to generate pseudo-labels, and multi-factor fusion and attention mechanisms are used to enhance the model’s robustness. Duan et al. [143] introduce a prompt-driven ZS-CMR method, which reduces the reliance on annotated triplets and textual data. By clustering images to generate pseudo-labels and using LLMs to create diverse textual modification instructions, this approach enables the model to perform composed retrieval using only image data.
In the second category, Jiang et al. [94] present a synthetic labeling framework that selects visually similar image pairs, generates differential captions using models like BLIP and LLMs, and filters outputs for semantic quality. Levy et al. [96] introduce LaSCo, a large-scale dataset formed by selecting image pairs that yield different answers to the same question, followed by GPT-3-generated transition text. LaSCo significantly enlarges the training set and improves model performance. Jang et al. [95] propose the Visual Delta Generator, which identifies suitable image pairs from auxiliary datasets and generates textual descriptions of visual differences to construct pseudo triplets. The CoLLM proposed by Huynh et al. [142] generates pseudo triplets by finding in-batch nearest neighbor images for each target in large-scale image-text pairs, interpolating reference features via Slerp, and creating modification texts with predefined templates and LLMs based on the captions. Noting that describing the differences between two images has been a long-standing research topic of interest [224, 225, 226, 227, 228].
5.1.2 Generating Language and Visual Content
Some works not only generate text but also generate images to further augment the triple data. Zhang et al. [97] propose CTI-IR, a GAN-based network that jointly trains generative and retrieval models to handle image generation and retrieval tasks. The model uses global-local collaborative discriminators to ensure semantic consistency between the generated image and the modification text by capturing both overall and fine-grained differences. Gu et al. [98] introduce a diffusion-based model for constructing a large-scale synthetic dataset called SynthTriplets18M. By leveraging existing image-caption datasets (e.g., COYO 700M, LAION-2B-en-aesthetic), they first generate diverse modified texts by replacing object terms in the original captions or editing operation via LLMs. Then, a latent diffusion model is used to generate high-quality target images based on the reference image and modification text. This approach enhances the model’s ability to handle diverse modifications and introduces capabilities like handling negative text and image mask conditions. Zhou et al. [73] propose one strategy, generating multi-modal triplets from large-scale image-text datasets. They use LLMs (GPT 3.5) to generate modification texts, which can be used to change certain objects, colors, positions, etc, and then employ stable diffusion to create target images, ensuring alignment between the modified text and the generated image.
Liu et al.  [111] propose an automated pipeline for a large-scale, synthetic, composed person retrieval dataset. First, a LLM generates textual quadruples that include reference and target image descriptions along with relative captions detailing appearance changes. Second, a fine-tuned generative model produces identity-consistent image pairs from these textual descriptions, with each pair split into reference and target images. Tu et al. [150] propose MRA-CIR, which generates pseudo-triplets by selecting semantically related image pairs and creating modification texts with a multimodal reasoning agent combining GNN-based visual reasoning, Transformer-based textual reasoning, and attention-based cross-modal fusion for effective compositional learning in ZS-CMR. Recently, Li et al. [149] propose an Imagine-and-Seek paradigm IP-CIR, where a diffusion model generates potential target image representations from reference images and modification texts, and these generated images serve as intermediate supervision to train the retrieval model, enhancing cross-modal alignment without real target images. Building on this, Wang et al. [148] design a generative framework that directly synthesizes target image representations from reference images and modification texts, jointly training a diffusion model with a contrastive matching module to enable effective retrieval without predefined triplets or pseudo-word tokens.
5.2 Noise/Uncertainty in Data
In addition to the volume of data, ensuring data quality is a critical focus in existing research. High-quality triplet data can significantly enhance model training effectiveness by mitigating the impact of noisy data, thereby improving retrieval performance. Data uncertainty and noise primarily manifest in two forms: (1) Content noise, where erroneous content is incorrectly deemed semantically relevant [94, 91, 98, 36, 92, 96]. For instance, the target image does not fully align with the retrieval intent specified by the modification text. This issue is particularly pronounced in automatically generated data. (2) Annotation uncertainty, where the target image in the triplet is not the sole positive example, as other images in the dataset may also satisfy the retrieval intent defined by the reference image and modification text [72, 99, 100, 101, 23, 32, 151]. Existing research has addressed these challenges through a series of studies. Below, we review and summarize these efforts in detail.
5.2.1 Content Noise
Existing methods primarily employ threshold-based filtering strategies to ensure semantic relevance and consistency. These strategies focus on measuring the similarity between reference images, modified texts, and target images to align with the intended modification intention. Jiang et al. [94] and Liu et al. [91] both utilize semantic similarity filtering to ensure consistency between generated modification texts and image pairs. Jiang et al. [94] screen modification texts within the language space, retaining triplets with high semantic similarity to both reference and target images. Similarly, Liu et al. [91] generate modification instructions using a LLM, followed by sentence transformation to produce edited descriptions. They apply a similarity-based filter to select triplets with high semantic relevance between edited descriptions and reference images. Gu et al. [98] design a CLIP-based filtering strategy, calculating CLIP similarity between reference and generated target images, as well as between original and modified texts, and applying thresholds to ensure visual and semantic consistency. They introduce directional CLIP similarity to evaluate alignment between image and text modifications and verify that generated images reflect specified keywords or instructions.
Lin et al. [36] introduce a similarity-based data augmentation method. Specifically, they utilize CLIP’s image encoder to extract features from all images in the training set, and then compute the cosine similarity between the reference or target images and other images in the training set. Based on the similarity scores, several of the most similar images are selected as new reference or target images, thereby generating pseudo-triplets. To ensure quality, the generated pseudo-triplets select the most similar images, constrained by predefined similarity thresholds and a limit on the number of pseudo-triplets per image to control quality and prevent overfitting. Hou et al. [92] introduce a random sampling strategy based on the 3-$\sigma$ rule to select challenging samples for fine-tuning. They generate pseudo-modification texts for unlabeled image pairs, calculate distance metrics, and define a candidate range using the 3-$\sigma$ rule. Randomly selected samples from this range mitigate noise while ensuring sufficient challenge, enhancing model robustness. Levy et al. [96] develop an analytical tool to detect redundant noise in queries by evaluating the contribution of different modalities. Queries with minimal impact from either image or text are identified as containing redundant information. By identifying and filtering out triplets that include such redundant information, the quality of the generated data can be enhanced. Liu et al.  [111] use a MLLM to evaluate and filter the generated triplets based on image quality, identity consistency, text-image alignment, and caption quality, retaining only high-quality examples that meet strict criteria.
5.2.2 Annotation Uncertainty
To address the issue of annotation uncertainty in triplets, where a single query may correspond to multiple valid target images, existing methods focus on mitigating false positives and enhancing semantic diversity. These approaches can be categorized into two main strategies: uncertainty-aware modeling and constructing evaluation datasets with multiple ground-truth labels.
Some works focus on modeling uncertainty in data. Wen et al. [72] introduce a matching regularization module based on target similarity distribution. It dynamically adjusts match scores by measuring similarities between the ground-truth target and candidate images within a mini-batch, allowing for partial matches rather than strictly treating all non-targets as negatives. Zhang et al. [99] propose the Consensus Network (Css-Net), which utilizes four diverse compositors to generate robust image-text embeddings. By applying Kullback-Leibler divergence loss, Css-Net encourages consistent outputs across compositors, mitigating individual biases. Chen et al. [100] present a rank-aware uncertainty framework with three modules: in-sample uncertainty models image features as Gaussian distributions; cross-sample uncertainty identifies shared positives across queries; and distribution regularization aligns source-target features. Chen et al. [101] further refine uncertainty modeling by introducing feature-space perturbations to simulate multi-granularity queries. Their model adjusts between one-to-one and one-to-many matching based on query precision via an uncertainty regularization module, thereby reducing false positives and enhancing retrieval accuracy.
Some studies have highlighted that uncertainty in data can become noise during model training, such as image-text pairs that are labeled as mismatched but are actually matched. To address this problem, researchers have proposed improved methods [229, 230, 231, 232, 233, 234, 235]. Huang et al. [234] propose Noisy Correspondence Rectifier, which divides data into two types, correct and noisy, based on the memory effect of neural networks, and then corrects the correspondence through an adaptive prediction model in a cooperative teaching manner. Biten et al. [229] propose a new strategy that defines a semantic adaptive boundary using the image captioning metric CIDEr and improves and optimizes it in the standard triple ranking loss. Yang et al. [235] proposed estimating soft labels for noisy annotations to reflect their true correspondence. Recently, Li et al. [151] propose pseudo-text enhancement for triple noisy annotations, which transforms the noise into a visual difference modeling problem, bridges the gap between real modifications and annotated text by generating adapters, and introduces learnable task-oriented prompts to replace reference images to construct independent queries and reduce the impact of visually irrelevant noise.
To better reflect real-world ambiguity, similar to the expansion of the dataset by supplementing the missing positive sample associations in the image-text pair [236], new datasets provide multiple valid targets per query. Baldrati et al. [23] create the CIRCO dataset, which annotates multiple relevant images for each query. They retrieve the top 100 candidates using retrieval models, filter to the 50 most visually similar images, and annotate all plausible matches, yielding an average of 4.53 ground truths per query. This reduces false negatives and improves evaluation robustness. Similarly, the CFQ dataset [32] offers multi-label annotations in the fashion domain. For each query, annotators label several target images from the same category as positive or negative based on their alignment with the textual description, providing a more comprehensive benchmark for fashion-oriented image retrieval.